Find the regression line of y on x for the following data:


X 40 70 50 60 80 50 90 40 60 60


Y 2.5 6.0 4.5 5.0 4.5 2.0 5.5 3.0 4.5 3.0


Draw the line. Estimate the value of y when x = 10.


1
Expert's answer
2021-02-03T03:52:49-0500
"\bar{x}=\dfrac{600}{10}=60, \bar{y}=\dfrac{40.5}{10}=4.05"


"SS_{xx}=\sum x_i^2-n\bar{x}^2=2400"

"SS_{yy}=\sum y_i^2-n\bar{y}^2=16.225"

"SS_{xy}=\sum x_iy_i-n\bar{x}\bar{y}=140"

"b=\dfrac{SS_{xy}}{SS_{xx}}=\dfrac{140}{2400}=\dfrac{7}{120}=0.0583"

"a=\bar{y}-b\bar{x}=4.05-\dfrac{7}{120}(60)=0.55"

"Y=a+bX"

"Y=0.55+0.0583X"


"r=\dfrac{SS_{xy}}{\sqrt{SS_{xx}}\sqrt{S_{yy}}}=\dfrac{140}{\sqrt{2400}\sqrt{16.225}}=0.7095"

"r^2=0.5033"

Moderate positive correlation.


"X=10"


"Y=0.55+0.0583(10)=1.133"